The point-slope form of an equation of a line is represented by y-y1=m(x-x1). What can be used to determine the slope?
Answer by checkley71(8403)

(Y-Y1)=m(X-X1) SHOULD BE (Y2-Y1)=m(X2-X1) OR (m)=(Y2-Y1)/(X2-X1)
NORMALLY THE SLOPE IS FOUND BY USING THE (X1,Y1) (X2,Y2) VALUES. USING THE FORMULA:
m=(Y2-Y1)/(X2-X1).
FOLLOWING IS AN EXAMPLE USING 2 POINTS.
(-3,4)(6,-2)
(-2-4)/(6+3)
-6/9=-2/3 IS THIS SLOPE.